# Semi-Supervised Risk Control via Prediction-Powered Inference

### v1

@@ II Background and Problem Setup
Denote by $D^{\text{cal}}_{L}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ a labeled calibration set such that $(X_{i},Y_{i})$ are sampled i.i.d. from $P_{XY}$ for all $1\leq i\leq n$. The feature $X$ takes values in $\mathcal{X}$ and the response $Y$ takes values in $\mathcal{Y}$. Let $\mathcal{T}_{q}(X)$ be a prediction of $X$ which depends on a pre-trained model $\hat{f}$ and is controlled by a hyper-parameter $q$. For example, in a classification task, $\mathcal{T}_{q}(X)$ is a prediction set that can be constructed by including all labels whose estimated class probabilities pass the threshold $1-q$. Finally, let $L(Y,\mathcal{T}_{q}(X))$ be a loss function that measures an error notion between the true response $Y$ and $\mathcal{T}_{q}(X)$; its expectation over future test samples $(X,Y)$, sampled i.i.d. from $P$ is the risk function, i.e., $R(q)=\mathbb{E}[L(Y,\mathcal{T}_{q}(X))]$. In the classification example, the loss could be the binary miscoverage loss $L(Y,\mathcal{T}_{q}(X))=\mathbbm{1}\{Y\notin\mathcal{T}_{q}(X)\}$, where $\mathbbm{1}\{A\}$ is the indicator function that equals one if $A$ occurs, and zero otherwise.
@@ III Related Work
In the context of conformal prediction, there have been attempts to overcome sample size barriers, however, all these attempts do not include using unlabeled data. In [27], the authors extend conformal prediction to include few-shot predictions with auxiliary tasks and create a meta-learning framework for constructing confident set-valued classifiers for new target tasks. Put simply, their approach leverages different but related tasks to enhance sample efficiency. This stands in contrast with our method which leverages unlabeled data. As such, one can view our method as an orthogonal direction to improve the sample efficiency of calibration methods. In [17], the authors focus on class-conditional coverage guarantees of conformal methods and suggest utilizing similarities between different classes to enhance the sample size of the calibration data for each class. Here, sample size issues become apparent when the number of classes is large, as the calibration set for each class becomes small. To mitigate this issue, the authors introduced a clustering approach, which combines samples from different classes that have similar prediction errors. In turn, this method attains a cluster-conditional rather than class-conditional coverage.
Lastly, independent work by [47], posted concurrently with our proposal, extends RCPS to sequential settings with an active labeling mechanism, ensuring valid risk control guarantees across all time steps. Similar to this paper, the authors of [47] also leverage PPI for variance reduction; however, their setup differs significantly from ours. First, the method in [47] focuses on an online setting, whereas we deal with an offline setup. Second, the validity of [47] only holds for monotonous risks, an assumption we do not make. In fact, it would be illuminating to further explore whether our method can extend their approach to handle non-monotonic risks.
@@ IV Warm Up: Non-Monotonic Risk Control via RCPS
Herein, we present a new result on the validity of RCPS, indicating that this framework in fact controls any general non-monotonic risk function and not merely monotonic ones, as argued in [11]. The procedure is similar to RCPS, however the pre-specified grid $\mathcal{Q}_{\text{grid}}$ might be different, especially in cases where we would not expect the risk to be small for larger values of $q$. Here, the choice of $\mathcal{Q}_{\text{grid}}$ is crucial as, by design, the procedure stops the first time $\hat{R}^{+}(q)$ does not fall below $\alpha$. As such, when the risk is not monotonous the grid $\mathcal{Q}_{\text{grid}}$ should be carefully designed in order to increase statistical efficiency [52, 13, 53]. With this design choice in mind, let $\mathcal{Q}_{\text{grid}}=\{q_{1},q_{2},\ldots,q_{M}\}$ be a pre-specified grid for the parameter $q$; in the monotonous case, $q_{1}=q_{\text{max}},q_{2}=q_{\text{max}}-\Delta,\ldots$ Then, the procedure proceeds by computing a UCB with error level $\delta$ at these specified points $(\hat{R}^{+}(q_{1}),\hat{R}^{+}(q_{1}),\ldots,\hat{R}^{+}(q_{M}))$. We iterate over this grid in order, from $q_{1}$ to $q_{M}$, and return the first value for which the UCB is not less than $\alpha$, as in (5).
@@
Under the assumption that the calibration and test set are i.i.d and given an upper confidence bound that satisfies (4), the RCPS procedure satisfies (2) for any general risk function (including non-monotonous risks).
This result indicates that the regular RCPS algorithm can be used to control any risk function in finite samples, for any data distribution, and regardless of the choice of the predictive model $\hat{f}$ used to construct $\mathcal{T}_{q}(X)$. This conclusion is not only of independent interest but also important for the derivation of the main contribution of this work: a semi-supervised calibration procedure with rigorous risk control.
Before describing this method, we pause to explain the high-level idea of the proof of Theorem 1. The proof is based on the learn then test framework presented in [13] which expands RCPS to control also non-monotonous risk functions. It reformulates risk control as a multiple-hypothesis testing task, testing for each $q\in\mathcal{Q}_{\text{grid}}$ whether the risk is not controlled. As such, the null hypothesis for a given $q$ is $H_{0}:\ R(q)>\alpha$. The idea here is to iterate over $\mathcal{Q}_{\text{grid}}$, calculate a p-value for each hypothesis, and stop the first time the null is not rejected. In turn, the validity of RSPC for non-monotonous risks follows from the duality between p-values and confidence bounds. The detailed algorithm and proof are provided in Appendix A-A.
@@ V Semi-Supervised RCPS: A General Loss
Next, we explain when and why this risk reduces the estimator’s variance compared to using only labeled data for constructing confidence intervals or upper confidence bounds. The variance of the empirical risk $\hat{R}_{L}(q)$ that uses only labeled data is $\hat{\sigma}^{2}_{L}={{\sigma}_{L_{i}}^{2}}/{n}$, where ${\sigma}_{L_{i}}^{2}$ is the variance of $L_{i}(q)$, which we cannot control. On the other hand, the variance of the empirical prediction-powered risk $\hat{R}_{\text{PP}}(q)$ in (6) is $\hat{\sigma}^{2}_{\text{PP}}=\hat{\sigma}^{2}_{U}+\hat{\sigma}^{2}_{\text{rect}}={{\sigma}^{2}_{\tilde{L}_{j}}}/{N}+{{\sigma}_{\tilde{L}_{i}-L_{i}}^{2}}/{n}$. Here, ${\sigma}^{2}_{\tilde{L}_{j}}$ is the variance of $\tilde{L}_{j}(q)$ and ${\sigma}_{\tilde{L}_{i}-L_{i}}^{2}$ is the variance of $\tilde{L}_{i}(q)-L_{i}(q)$. Notably, the variance of $\hat{R}_{\text{PP}}(q)$ is the sum of these two variances, since $\{\tilde{L}_{i}(q)-L_{i}(q)\}_{i=1}^{n}$ and $\{\tilde{L}_{j}(q)\}_{j=n+1}^{n+N}$ are evaluated on independent samples. With a large number $N$ of unlabeled samples, ${\hat{\sigma}^{2}_{\tilde{L}_{j}}}/{N}$ is close to zero. Moreover, $\hat{\sigma}_{\tilde{L}_{i}-L_{i}}^{2}$ becomes smaller as the imputation accuracy increases. In turn, when the imputed labels are reasonably accurate, $\hat{\sigma}_{\text{PP}}^{2}$ is anticipated to be smaller than $\hat{\sigma}^{2}_{L}$. Another way to view the benefit of PPI is this: when the imputations are accurate, $\hat{R}_{\text{rect}}(q)$ in (6) is approximately zero, while $\hat{R}_{U}(q)$ is close to the true risk as the sample size $N$ is large. This, in turn, can be beneficial to obtain a tighter UCB, as we show next. Nonetheless, even with inaccurate imputations, the UCB will remain valid as $\hat{R}_{\text{PP}}(q)$ is unbiased, though we would not gain improvement in statistical efficiency compared to using only labeled data.
@@
In practice, we recommend using the state-of-the-art WSR UCB described in Algorithm 4 of Appendix A-B to attain finite-sample validity or build upon the central limit theorem (CLT) to construct an asymptotically valid UCB. The latter can be less conservative, at the cost of relying on stronger assumptions. Notably, the random variables $\{W_{i}\}_{i=1}^{n}$ have different support than the original losses $L_{i}(q)$. For example, in the binary case, the loss is bounded between 00 and $1$ while $W_{i}$ is bounded between $-1$ and $2$. Therefore, one should account for this discrepancy when constructing the bound, especially because various concentration inequalities, such as WSR, assume the variables are bounded between 00 and $1$. To tackle this, in Algorithm 4 of Appendix A-B we present a simple extension of the original WSR for losses bounded between any two numbers $A$ and $B$, similar to [54].
@@
Note that when resorting to asymptotic guarantees, one can further enhance the statistical efficiency of PPI, as described in [43]. The idea is to introduce a hyper-parameter $\lambda\in\mathbb{R}$ that controls the reliance on the unlabeled data. This is done by multiplying by $\lambda$ the imputed losses $\frac{1}{N}\sum_{j=1}^{N}\tilde{L}_{j}(q)-\frac{1}{n}\sum_{i=1}^{n}\tilde{L}_{i}(q)$ in (6). This hyper-parameter $\lambda$ can be set in a data-driven fashion, where the higher the accuracy the larger the $\lambda$ will be. Nonetheless, $\lambda$ cannot be tuned when providing finite-sample guarantees without assuming access to additional hold-out data. For this reason, in our method, we set $\lambda=1$ by default.
@@ VI Specialized Semi-Supervised RCPS: a Binary Loss
where $R^{\text{clip}}_{\text{rect}}(q)=\mathbb{E}[(L(q)-\tilde{L}(q))_{+}].$ The operator $(z)_{+}$ returns $\max\{z,0\}$. While $(L(q)-\tilde{L}(q))$ can have values in $\{-1,0,1\}$, its clipped version $(L(q)-\tilde{L}(q))_{+}$ can only achieve values in $\{0,1\}$. In turn, the clipping enables us to use the exact Clopper–Pearson UCB for both $R^{\text{clip}}_{\text{rect}}(q)$ and $R_{U}(q)$, as the latter involves a binary loss by definition.
@@ VII-A Few Shot Calibration: Image Classification
To formalize the prediction set function, we follow the conformal prediction approach and define $s(X,Y)\in\mathbb{R}$ as a real-valued score function that represents the degree of agreement between the model’s prediction $\hat{f}(X)$ and the true label $Y$. This score function can take any form, but we use the convention that a higher score indicates a higher prediction error. With this in place, we define the prediction sets as $\mathcal{T}_{q}(X)=\{y\in\mathcal{Y}:s(X,y)\leq q\}$, i.e., we include in $\mathcal{T}_{q}(X)$ all the labels $y$ whose corresponding scores are smaller than or equal to $q$. In the following experiments, we use the adaptive prediction sets (APS) score, introduced in [56], which sums the sorted estimated class probabilities derived by the classifier $\hat{f}(X)$ up to the rank of $Y$. The exact formulation of this score function is given in Appendix C-A.
@@
  • RCPS imp: a naive, invalid RCPS method, applied to a calibration set that includes the hold-out labeled data as well as the imputed unlabeled samples. We refer to this as a naive approach as it does not account for erroneous imputations. We include this method in our experiment to underscore the need for our rigorous semi-supervised methods that leverage unlabeled data while accounting for incorrect imputations to ensure valid results. In our experiments, we employ a slightly modified version of APS for this method, which is even more conservative than the standard APS in terms of coverage, see Appendix C-A for details.
@@
To implement the semi-supervised calibration methods, we impute the unknown labels of the unlabeled samples using the most likely, top-1 predictions obtained from a pre-trained Resnet50 classifier, which has a top-1 accuracy of approximately 81%. This classifier is also used to compute the APS scores and construct the prediction sets. Consequently, by design, we have $\tilde{L}(q)\leq L(q)$ almost surely, since $s(X,\tilde{Y})\leq s(X,Y)$ almost surely. This is an example of a situation where the clipping operator does not introduce any bias as $R_{\text{PP}}(q)=R_{\text{PP}}^{\text{clip}}(q)$ from (17).
@@
To create the test set as well as the labeled and unlabeled calibration sets, we randomly split the 50,000 samples from the original ImageNet validation set. Specifically, 25,000 samples are used for testing, and the remaining 25,000 samples are used to form the two disjoint labeled and unlabeled calibration sets. In each experiment, we evaluate the empirical coverage rate—defined as one minus the empirical risk—and the prediction set sizes obtained by each calibration method. We set the error levels of the calibration methods to $(\alpha,\delta)=(0.15,0.1)$. For the SS-RCPS BIN, we fix $(\delta_{1},\delta_{2})=(0.01,0.09)$ when the labeled calibration set is of size less than 1000, and $(\delta_{1},\delta_{2})=(0.05,0.05)$ otherwise.
@@
Figure 1: ImageNet classification experiment: the effect of labeled calibration set size. The top panel shows the coverage rate of the constructed prediction sets, while the bottom panel displays the average size of the prediction sets. Both performance metrics are presented as a function of the number of labeled calibration samples and evaluated over 100 different random splits of the data. Refer to the main text for a description of the five different methods being compared.
@@
Figure 1 displays the empirical distribution of the coverage and set-size metrics as a function of the number of labeled calibration samples. Following the top panel of Figure 1, we can see that the vanilla RCPS lab controls the risk with high probability, as expected. Observe that when the sample size of the labeled calibration set is small the empirical coverage rate is conservatively controlled and the coverage variance is relatively large. In contrast, the semi-supervised calibration with the CLT bound (SS-RCPS CLT) is less conservative in terms of coverage, but this bound is not always valid. For example, with 130 labeled samples, the coverage rate is below the desired 85% nominal level 18% of the time instead of 10%. The finite-sample semi-supervised approach for general loss with the WSR bound (SS-RCPS WSR) is valid across the board, yet, this approach is more conservative than SS-RCPS CLT, especially for small sample sizes. Notably, SS-RCPS WSR tends to surpass both the vanilla RCPS lab and the specialized semi-supervised calibration method (SS-RCPS BIN) as the sample size increases. The specialized method, however, outperforms both the vanilla RCPS lab and SS-RCPS WSR in the small sample-size regime. Lastly, the naive RCPS imp method that utilizes all imputed labels without correction leads to an invalid coverage rate. This is anticipated as $\tilde{L}(q)\leq L(q)\Rightarrow R_{U}(q)\leq R(q)$ in our experiments. In turn, by naively augmenting the small labeled data with the large imputed data we control a risk that is smaller than the desired $R(q)$. Consequently, the naive calibrated parameter $\hat{q}$ is too small, failing to achieve the desired coverage rate. We now turn to the bottom panel of Figure 1, portraying the average size of the prediction sets. We can see that the methods that attain lower coverage levels result in smaller prediction sets, as expected.
@@
In the next experiment, we examine how the accuracy of imputed labels affects the performance of our proposed semi-supervised calibration method. We fix the size of the labeled calibration set at 130 and synthetically increase the imputation accuracy by gradually replacing the incorrect imputed labels obtained by the classifier with the true labels. We focus on the SS-RCPS BIN method, which—as shown in Figure 1—performs best in this challenging small sample-size regime. Importantly, similarly to the experiment in Figure 1, we construct the prediction sets using the same pre-trained Resnet classifier. The results, summarized in Figure 2, show that as imputation accuracy improves, the coverage variance decreases and the gap between the average coverage and the nominal level narrows. In line with this trend, the average size of prediction sets decreases as the accuracy of the imputed labels increases. Further, we can see that the average coverage remains somewhat conservative even with accurate imputations due to the budget splitting of the specialized method. All in all, this experiment highlights that significant improvement in calibration performance can be obtained when relying on accurate imputed labels, as expected. At the same time, it is crucial to note that the risk-controlling guarantee always holds, regardless of the imputation accuracy.

### v2

@@ II Background and Problem Setup
[…]
@@ III Related Work
In the context of conformal prediction, there have been attempts to overcome sample size barriers, however, all these attempts do not include using unlabeled data. In [27], the authors extend conformal prediction to include few-shot predictions with auxiliary tasks and create a meta-learning framework for constructing confident set-valued classifiers for new target tasks. Put simply, their approach leverages different but related tasks to enhance sample efficiency. This stands in contrast with our method which leverages unlabeled data. As such, one can view our method as an orthogonal direction to improve the sample efficiency of calibration methods. In [17], the authors focus on class-conditional coverage guarantees of conformal methods and suggest utilizing similarities between different classes to enhance the sample size of the calibration data for each class. Here, sample size issues become apparent when the number of classes is large, as the calibration set for each class becomes small. To mitigate this issue, the authors introduced a clustering approach, which combines samples from different classes that have similar prediction errors. In turn, this method attains a cluster-conditional rather than class-conditional coverage. Other approaches to improve statistical efficiency include cross-validation [47], meta-learning [48, 49], and more efficient use of data for model- and score-selection [50].
Lastly, independent work by [51], posted concurrently with our proposal, extends RCPS to sequential settings with an active labeling mechanism, ensuring valid risk control guarantees across all time steps. Similar to this paper, the authors of [51] also leverage PPI for variance reduction; however, their setup differs significantly from ours. First, the method in [51] aims to reduce variance per input to improve the effectiveness of active labeling, but this approach would have no impact in the offline setting we consider in this paper, as it does not involve active labeling. Second, the validity of [51] only holds for monotonous risks, an assumption we do not make. In fact, it would be illuminating to further explore whether our method can extend their approach to handle non-monotonic risks.
@@ IV Warm Up: Non-Monotonic Risk Control via RCPS
Herein, we present a new result on the validity of RCPS, indicating that this framework in fact controls any general non-monotonic risk function and not merely monotonic ones, as argued in [11]. The procedure is similar to RCPS, however the pre-specified grid $\mathcal{Q}_{\text{grid}}$ might differ, especially when we would not expect the risk to be small for larger $q$ values. Here, the choice of $\mathcal{Q}_{\text{grid}}$ is crucial as, by design, the procedure stops the first time $\hat{R}^{+}(q)$ is not below $\alpha$. As such, when the risk is not monotonous, the grid $\mathcal{Q}_{\text{grid}}$ should be carefully designed to increase statistical efficiency [56, 13, 57]. For example, our early time series classification experiment in Section VII-B involves tuning a complex, high-dimensional hyperparameter vector. To do so, we adopt the approach proposed in [56], which is related to the methods introduced in [57, 58]. With this design choice in mind, let $\mathcal{Q}_{\text{grid}}=\{q_{1},q_{2},\ldots,q_{M}\}$ be a pre-specified grid for the parameter $q$; in the monotonous case, $q_{1}=q_{\text{max}},q_{2}=q_{\text{max}}-\Delta,\ldots$ Then, the procedure proceeds by computing a UCB with error level $\delta$ at these specified points $(\hat{R}^{+}(q_{1}),\hat{R}^{+}(q_{2}),\ldots,\hat{R}^{+}(q_{M}))$. We iterate over this grid in order, from $q_{1}$ to $q_{M}$, and return the first value for which the UCB is not less than $\alpha$, as in (5).
@@
Assuming the calibration and test set are i.i.d. and given an upper confidence bound that satisfies (4), the RCPS procedure satisfies (2) for any general risk function (including non-monotonous risks).
This result indicates that the regular RCPS algorithm can be used to control any risk function in finite samples, for any data distribution, and regardless of the choice of the predictive model $\hat{f}$ used to construct $\mathcal{T}_{q}(X)$. This conclusion is not only of independent interest, but also important for the derivation of the main contribution of this work: a semi-supervised calibration procedure with rigorous risk control.
We pause to explain the high-level idea of the proof of Theorem 1. It is based on the learn then test framework presented in [13] which expands RCPS to control also non-monotonous risk functions. It reformulates risk control as a multiple-hypothesis testing task, testing for each $q\in\mathcal{Q}_{\text{grid}}$ whether the risk is not controlled. As such, the null hypothesis for a given $q$ is $H_{0}:\ R(q)>\alpha$. The idea is to iterate over $\mathcal{Q}_{\text{grid}}$, calculate a p-value for each hypothesis, and stop the first time the null is not rejected. We show that RCPS and learn then test are equivalent when implemented with a search grid $\mathcal{Q}_{\text{grid}}=\{q_{1},q_{2},...,q_{M}\}$ of the same ordering. This is due to the duality between p-values and confidence bounds. As a result, the validity of RCPS for non-monotonous risks follows naturally. The detailed algorithm and proof are provided in Appendix A-A.
@@ V Semi-Supervised RCPS: A General Loss
Next, we explain when and why this risk reduces the estimator’s variance compared to using only labeled data for constructing confidence intervals or upper confidence bounds. The variance of the empirical risk $\hat{R}_{L}(q)$ that uses only labeled data is $\hat{\sigma}^{2}_{L}={{\sigma}_{L_{i}}^{2}}/{n}$, where ${\sigma}_{L_{i}}^{2}$ is the variance of $L_{i}(q)$, which we cannot control. On the other hand, the variance of the empirical prediction-powered risk $\hat{R}_{\text{PP}}(q)$ in (6) is $\hat{\sigma}^{2}_{\text{PP}}=\hat{\sigma}^{2}_{U}+\hat{\sigma}^{2}_{\text{rect}}={{\sigma}^{2}_{\tilde{L}_{j}}}/{N}+{{\sigma}_{\tilde{L}_{i}-L_{i}}^{2}}/{n}$. Here, ${\sigma}^{2}_{\tilde{L}_{j}}$ is the variance of $\tilde{L}_{j}(q)$ and ${\sigma}_{\tilde{L}_{i}-L_{i}}^{2}$ is the variance of $\tilde{L}_{i}(q)-L_{i}(q)$. Notably, the variance of $\hat{R}_{\text{PP}}(q)$ is the sum of these two variances, since $\{\tilde{L}_{i}(q)-L_{i}(q)\}_{i=1}^{n}$ and $\{\tilde{L}_{j}(q)\}_{j=n+1}^{n+N}$ are evaluated on independent samples. With a large number $N$ of unlabeled samples, ${\sigma^{2}_{\tilde{L}_{j}}}/{N}$ is close to zero. Moreover, $\sigma_{\tilde{L}_{i}-L_{i}}^{2}$ becomes smaller as the imputation accuracy increases. In turn, when the imputed labels are reasonably accurate, $\hat{\sigma}_{\text{PP}}^{2}$ is anticipated to be smaller than $\hat{\sigma}^{2}_{L}$. Another way to view the benefit of PPI is this: when the imputations are accurate, $\hat{R}_{\text{rect}}(q)$ in (6) is approximately zero, while $\hat{R}_{U}(q)$ is close to the true risk as the sample size $N$ is large. This, in turn, can be beneficial to obtain a tighter UCB, as we show next. Nonetheless, even with inaccurate imputations, the UCB will remain valid as $\hat{R}_{\text{PP}}(q)$ is unbiased, though we would not gain improvement in variance reduction compared to using only labeled data, and might even lose statistical efficiency. At the end of this section and in Appendix C-A2, we discuss a method that follows [43] and mitigates this negative effect by introducing a hyperparameter to control the influence of unlabeled data, thereby enabling more controlled variance reduction in the PPI risk estimator.
@@
In practice, we recommend using the state-of-the-art WSR UCB described in Algorithm 4 of Appendix A-B to attain finite-sample validity or build upon the central limit theorem (CLT) to construct an asymptotically valid UCB. The latter can be less conservative, at the cost of relying on stronger assumptions. Notably, the random variables $\{W_{i}\}_{i=1}^{n}$ have different support from the original losses $L_{i}(q)$. For example, in the binary case, $L_{i}(q)\in[0,1],$ while $W_{i}\in[-1,2]$. One should account for this discrepancy when constructing the UCB, especially because various concentration inequalities, such as WSR, assume the variables are bounded between 0 and 111. To tackle this, in Algorithm 4 of Appendix A-B we present a simple extension of the original WSR for losses bounded between any two numbers $A$ and $B$, similar to [59]. Note that linearly scaling the loss to the [0,1][0,1][ 0 , 1 ] range is also possible. As demonstrated in Appendix A-B, such a linear scaling approach achieves comparable power to the generalized WSR UCB method, which extends beyond the standard [0,1][0,1][ 0 , 1 ] loss.
@@
Note that the UCB constructed with the CLT is given by $\bar{W}+z_{1-\delta}\cdot\sqrt{\frac{\sigma^{2}_{W}}{n}}$ where $\bar{W}=\frac{1}{n}\sum_{i=1}^{n}W_{i}$, $z_{1-\delta}$ is the $1-\delta$ quantile of the standard normal distribution, and $\sigma^{2}_{W}$ is the variance of $W_{i}$. Since we assume the labeled and unlabeled points are i.i.d., the above CLT-based UCB is equivalent to a one-sided confidence interval constructed with Algorithm 1 of [34]. Thus, when resorting to asymptotic guarantees, one can further enhance the statistical efficiency of PPI, as described in [43]. The idea is to introduce a hyper-parameter $\lambda\in\mathbb{R}$ that controls the reliance on the unlabeled data. This is done by multiplying by $\lambda$ the imputed losses $\frac{1}{N}\sum_{j=n+1}^{n+N}\tilde{L}_{j}(q)-\frac{1}{n}\sum_{i=1}^{n}\tilde{L}_{i}(q)$ in (6). This hyper-parameter $\lambda$ can be set in a data-driven fashion, where the higher the accuracy the larger the $\lambda$ will be. Nonetheless, $\lambda$ cannot be tuned when providing finite-sample guarantees without assuming access to additional hold-out data. For this reason, in our method, we set $\lambda=1$ by default. See Appendix C-A2 for further details and additional baselines involving this parameter.
@@ VI Specialized Semi-Supervised RCPS: a Binary Loss
[…]
@@ VII-A Few Shot Calibration: Image Classification
To formalize the prediction set function, we follow the conformal prediction approach and define $s(X,Y)\in\mathbb{R}$ as a real-valued score function that represents the degree of agreement between the model’s prediction $\hat{f}(X)$ and the true label $Y$. This score function can take any form, but we use the convention that a higher score indicates a higher prediction error. With this in place, we define the prediction sets as $\mathcal{T}_{q}(X)=\{y\in\mathcal{Y}:s(X,y)\leq q\}$, i.e., we include in $\mathcal{T}_{q}(X)$ all the labels $y$ whose corresponding scores are smaller than or equal to $q$. In the following experiments, we use the adaptive prediction sets (APS) score, introduced in [61], which sums the sorted estimated class probabilities derived by the classifier $\hat{f}(X)$ up to the rank of $Y$. The exact formulation of this score function is given in Appendix C-A1.
@@
  • RCPS imp: a naive, invalid RCPS method, applied to a calibration set that includes the hold-out labeled data as well as the imputed unlabeled samples. We refer to this as a naive approach as it does not account for erroneous imputations. We include this method in our experiment to underscore the need for our rigorous semi-supervised methods that leverage unlabeled data while accounting for incorrect imputations to ensure valid results. In our experiments, we employ a slightly modified version of APS for this method, which is even more conservative than the standard APS in terms of coverage, see Appendix C-A1 for details.
@@
To implement the semi-supervised calibration methods, we impute the unknown labels of the unlabeled samples using the most likely, top-1 predictions obtained from a pre-trained Resnet50 classifier, which has a top-1 accuracy of approximately 81%. This classifier is also used to compute the APS scores and construct the prediction sets. Consequently, by design, we have $\tilde{L}(q)\leq L(q)$ almost surely, since $s(X,\tilde{Y})\leq s(X,Y)$ almost surely. This is an example of a situation where the clipping operator does not introduce any bias, as $R_{\text{PP}}(q)=R_{\text{PP}}^{\text{clip}}(q)$ from (17).
@@
To create the test set as well as the labeled and unlabeled calibration sets, we randomly split the 50,000 samples from the original ImageNet validation set. Specifically, 25,000 samples are used for testing, and the remaining 25,000 samples are used to form the two disjoint labeled and unlabeled calibration sets. In each experiment, we evaluate the empirical coverage rate—defined as one minus the empirical risk—and the prediction set sizes obtained by each calibration method. We set the error levels to $(\alpha,\delta)=(0.15,0.1)$. For the SS-RCPS BIN, we fix $(\delta_{1},\delta_{2})=(0.01,0.09)$ when the labeled calibration set is of size less than 1000, and $(\delta_{1},\delta_{2})=(0.05,0.05)$ otherwise.
@@
Figure 1: ImageNet classification experiment: the effect of labeled calibration set size. The top panel shows the coverage rate of the constructed prediction sets, while the bottom panel displays the average size of the prediction sets. Both performance metrics are presented as a function of the number of labeled calibration samples and evaluated over 100 different random splits of the data. Refer to the main text for a description of the five different methods compared.
@@
Figure 1 displays the empirical distribution of the coverage and set-size metrics as a function of the number of labeled calibration samples. Following the top panel of Figure 1, we can see that the vanilla RCPS lab controls the risk with high probability, as expected. Observe that when the sample size of the labeled calibration set is small the empirical coverage rate is conservatively controlled and the coverage variance is relatively large. In contrast, the semi-supervised calibration with the CLT bound (SS-RCPS CLT) is less conservative in terms of coverage, but this bound is not always valid. For example, with 130 labeled samples, the coverage rate is below the desired 85% nominal level 18% of the time instead of 10%. The finite-sample semi-supervised approach for general loss with the WSR bound (SS-RCPS WSR) is valid across the board, yet, this approach is more conservative than SS-RCPS CLT, especially for small sample sizes. Notably, SS-RCPS WSR tends to surpass both the vanilla RCPS lab and the specialized semi-supervised calibration method (SS-RCPS BIN) as the sample size increases. The specialized method, however, outperforms both the vanilla RCPS lab and SS-RCPS WSR in the small sample-size regime. Lastly, the naive RCPS imp method that utilizes all imputed labels without correction leads to an invalid coverage rate. This is anticipated as $\tilde{L}(q)\leq L(q)\Rightarrow R_{U}(q)\leq R(q)$ in our experiments. In turn, by naively augmenting the small labeled data with the large imputed data we control a risk that is smaller than the desired $R(q)$. Consequently, the naive calibrated parameter $\hat{q}$ is too small, failing to achieve the desired coverage rate. Finally, the bottom panel of Figure 1 portrays the average size of the prediction sets. As expected, methods that achieve lower coverage levels result in smaller prediction sets.
@@
In the next experiment, we examine how the accuracy of imputed labels affects the performance of our semi-supervised calibration method. We fix the size of the labeled calibration set at 130 and synthetically increase the imputation accuracy by gradually replacing the incorrect imputed labels obtained by the classifier with the true labels. We focus on the SS-RCPS BIN method, which—as shown in Figure 1—performs best in this challenging small sample-size regime. Similarly to the experiment in Figure 1, we construct the prediction sets using the same pre-trained Resnet classifier. The results, summarized in Figure 2, show that as imputation accuracy improves, the coverage variance decreases and the gap between the average coverage and the nominal level narrows. In line with this trend, the average size of prediction sets decreases as the accuracy of the imputed labels increases. Further, we can see that the average coverage remains somewhat conservative even with accurate imputations due to the budget splitting of the specialized method. Overall, this experiment highlights that significant improvement in calibration performance can be obtained when relying on accurate imputed labels, as expected. Importantly, the risk-controlling guarantee holds, regardless of the imputation accuracy.